OpenStudy (anonymous):

how to do these surd equations? x-1 = square root 5x-9 x-3 = square root 30 - 2x

OpenStudy (anonymous):

\[x-1 = \sqrt{5x-9}\]

OpenStudy (anonymous):

and \[x-3 = \sqrt{30-2x}\]

OpenStudy (anonymous):

the question is to solve for x and y ?

OpenStudy (anonymous):

No, just for x

terenzreignz (terenzreignz):

Well... the seemingly natural way to do it is to square both sides of the equation...

terenzreignz (terenzreignz):

Take the first one, for instance... \[\Large (x-1)^2 = \left(\sqrt{5x-9}\right)^2\] Squaring would cancel the radical, leaving you with... \[\Large x^2 -2x + 1 = 5x -9\]

OpenStudy (anonymous):

for the first one x=5 ans x=2

OpenStudy (anonymous):

you just have to square both sides first then factorize the square

OpenStudy (anonymous):

and for the second one x=7 and x=-3

OpenStudy (anonymous):

did that make sense ?

OpenStudy (anonymous):

is there meant to be two solutions for those questions?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

cuz when square both side you will get x^2 in the left side and thar tells there are two independent solutions for the equation

OpenStudy (anonymous):

can you show me your working because i got a completely different answer

OpenStudy (anonymous):

sure just a sec

OpenStudy (anonymous):